Solve the following ode using Laplace transform: y - 5y = f(t); y(0) - 1 t; Ost<1 f(t) = 0; t21
Answer #1

Boit?! Given y-5y = f(t); y 10) = da) = It; osta + f(t) = t[uct) - u(t-1)] + b)ult -J = tulti - tult-1) = tults - (t-1+1) ulApplying partial fractions confort 12+1 123-5) A + B 32 + C S 3-5 S²+1 3215-9) ASIS-5) + B (5-5)+C52 S²15-5) 52 (A+C 1.52 +5ABy partial fractions -f A 3 + 543-5 + 3-5 -S 5?3-5) As Is-5) + B15-5) +C52 1 S215-5) & 1-5= (A+C) 52+(-5A +B)s - 5B Comp arin1-S 4 52(5-5) 255 4 2515-5 2 552 .: YO3) = [ Yes -425 f saya 26/25 t [45 / j2 4/25 Als Applying Inverse Laplace Transform YHAnsurer is y(t) = 1 / 3 t/m t 5 t 26 25 est -( -- 25 Ilu 5 $ 2542-1)uto ult=1)